Write the equation of a polynomial function in standard form that has zeros at -3, -2 and -2.
(X+ )(X+ )(X+ )=

Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

Write the equation of a polynomial function in standard form that has zeros at -3, -2 and -2.
(X+3)(X+2)(X+2)
------
= (x+3)(x^2 + 4x + 4)
-----
= x^3 + 4x^2 + 4x + 3x^2 + 12x + 12
-----
= x^3 + 7x^2 + 16x + 12
------------
Cheers,
Stan H.
-------------

Answer by ikleyn(52879)

.
Write the equation of a polynomial function in standard form that has zeros at -3, -2 and -2.
-----------------------------------------------

(X+3)(X+2)(X+2)= . . .

Calculate it yourself if you want or if you need.